OpenStudy (anonymous):

Using the appropriate Double Angle Formula, simplify the expression: 2(sin25°)(cos25°)

OpenStudy (raden):

use the identity : 2sin(x)cos(x) = sin(2x)

OpenStudy (anonymous):

The answer book says the answer is sin(50°). How?

OpenStudy (raden):

if u subtitute x=25 to formula above, giving us 2sin(25)cos(25) = sin(2*25) = sin(50), right ?